Jim is twice as old as Karen. Eight years from now, Jim will be twenty years older than Karen. How old is Jim now?
Answer by jorel1380(2518)

Let J=Jim, K=Karen;
J=2K
J+8=K+8+20
2K+8=K+28
K=20
J=40